1+(1/2)square+(1/3)square+.....infinity=(pie)square upon 6

Grade:11

1 Answers

Aman Bansal
12 years ago

Dear Rajesh,

Let f(x) = x^2   This is an even function of x so we know that NONE of 
the sine terms of the Fourier series can be present. We need only look 
at the a(i) terms.

For all a(n)  we get pi.a(n) = INT(-pi to pi)[x^2.cos(nx)dx]

When  n = 0 this gives    a(0) = (2/3)pi^2

When n greater than 0 integration by parts in 3 steps gives

                      a(n) = 4(-1)^n/n^2

The Fourier series is  pi^2/3 + 4.SUM(1 to infinity)(-1)^n/n^2 cos(nx)

With x=pi, we get

      pi^2 = pi^2/3 + 4.SUM(-1)^n/n^2 cos(n.pi)

and since cos(n.pi)= (-1)^n  this produces + s for each term

      2.pi^2/3  =  4.SUM[1 + 1/2^2 + 1/3^2 + ..... to infinity]

         pi^2/6 =  1 + 1/2^2 + 1/3^2 + 1/4^2 + ..... to infinity

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert